OpenStudy (anonymous):

Simplify the expression 17x-24x+13x. Your answer should have the variable x in it only once.

OpenStudy (solomonzelman):

17 - 24 - 13 = what ? Then 17x - 24x - 13x = the same thing, but with an ` x` in front of it

OpenStudy (solomonzelman):

I mean with an x after it.

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

can you tell me what you got ?

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

You see, if the expression is not 17x-24x+13x, but is 17*2-24*2+13*2, the answer of it will be (17-24+13)*2=6*2=12. Now just change "2" into "x". Then write it in the form of (17-24+13)*2.

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

yeah I GOT IRIGHT

OpenStudy (anonymous):

Great!